A galvanic cell is constructed by coupling Ag and Cd electrodes dipped in 0.5 M Silver Nitrate and 0.25 M CdS04 respectively at 25°C. Write the cell scheme, cel! reaction, and calculate emf of the cell. Given that SRPs of Ag and Cd are +0.80 V and ▪ -0.40 V respectively.


1
Expert's answer
2021-01-11T03:56:45-0500

Cell scheme: Cd2+ (0.25M)/Cd//Ag+ (0.5M)/Ag

Cell reaction: 2Ag+ + Cd----> Ag + Cd2+

E°= E°Ag - E°Cd

E°= 0.80 - (-0.40)

E°= 1.20V


Now let's find E using Nernst equation

E=E° - 0.059/n log[Cd2+]/[Ag+

E= 1.20 - 0.0591/2 log[0.25]/[0.50]²

E= 1.20- 0.02955log1

E= 1.20 - 0

E= 1.20V